# Generalized Scattering Matrix of Antenna: Moment Solution, Compression Storage and Application

### v3

@@ II-A Magnetic-type Equations—Magnetic Currents and PEC Boundary
To discretize (13) using MoM, we expand $\bm{J}^{e}=\bm{J}^{e}_{\mathrm{A}}\cup\bm{J}^{e}_{\mathrm{P}}$ and $\mathrm{j}\bm{J}^{m}$ through RWG basis functions $\left\{\bm{\psi}_{i}\right\}$:
$\bm{J}^{e}=\sum_{i}{I_{i}^{e}\bm{\psi}_{i},\mathrm{j}\bm{J}^{m}}=\sum_{i}{I_{i}^{m}\bm{\psi}_{i}}.$ (14)
@@
$\begin{split}&\mathbf{Z}^{\mathrm{M}}=\begin{bmatrix}\mathrm{j}k\eta_{0}\mathbf{L}&-\mathrm{j}\mathbf{K}^{+}\\ -\mathrm{j}\mathbf{K}^{-}&\mathrm{j}k/\eta_{0}\mathbf{L}+\mathbf{G}^{\mathrm{M}}\\ \end{bmatrix}\\ &\mathbf{I}=\begin{bmatrix}\mathbf{I}^{e}\\ \mathbf{I}^{m}\\ \end{bmatrix},\mathbf{V}^{\mathrm{inc}}=\begin{bmatrix}\mathbf{V}^{e}\\ \mathrm{j}\mathbf{V}^{m}\\ \end{bmatrix}=\begin{bmatrix}\left<\left\{\bm{\psi}_{i}\right\},\bm{E}^{\mathrm{inc}}\right>_{\mathrm{A}}\\ \mathrm{j}\left<\left\{\bm{\psi}_{i}\right\},\bm{H}^{\mathrm{inc}}\right>_{\mathrm{P}}\\ \end{bmatrix}\\ &\mathbf{V}^{\mathrm{M}}=\begin{bmatrix}\mathbf{O}\\ \mathrm{j}\mathbf{V}_{\mathrm{P}}^{m}\\ \end{bmatrix},\left[\mathbf{V}_{\mathrm{P}}^{m}\right]_{i}=-2\left<\bm{\psi}_{i},\sum_{\alpha}{v_{\alpha}\frac{\hat{n}\times\mathbf{e}_{\alpha}}{\sqrt{\eta_{\alpha}}}}\right>_{\mathrm{P}}\end{split}$ (16)
@@ V-A Validation example
Figure 5: Gain pattern at 3.5 GHz. (a) E-plane pattern, (b) H-plane pattern.

### v4

@@ II-A Magnetic-type Equations—Magnetic Currents and PEC Boundary
To discretize (13) using MoM, we expand $\bm{J}^{e}=\bm{J}^{e}_{\mathrm{A}}\cup\bm{J}^{e}_{\mathrm{P}}$ and $\bm{J}^{m}$ through RWG basis functions $\left\{\bm{\psi}_{i}\right\}$:
$\bm{J}^{e}=\sum_{i}{I_{i}^{e}\bm{\psi}_{i},\bm{J}^{m}}=\sum_{i}{I_{i}^{m}\bm{\psi}_{i}}.$ (14)
@@
$\begin{split}&\mathbf{Z}^{\mathrm{M}}=\begin{bmatrix}\mathrm{j}k\eta_{0}\mathbf{L}&-\mathrm{j}\mathbf{K}^{+}\\ -\mathrm{j}\mathbf{K}^{-}&\mathrm{j}k/\eta_{0}\mathbf{L}+\mathbf{G}^{\mathrm{M}}\\ \end{bmatrix}\\ &\mathbf{I}=\begin{bmatrix}\mathbf{I}^{e}\\ \mathrm{j}\mathbf{I}^{m}\\ \end{bmatrix},\mathbf{V}^{\mathrm{inc}}=\begin{bmatrix}\mathbf{V}^{e}\\ \mathrm{j}\mathbf{V}^{m}\\ \end{bmatrix}=\begin{bmatrix}\left<\left\{\bm{\psi}_{i}\right\},\bm{E}^{\mathrm{inc}}\right>_{\mathrm{A}}\\ \mathrm{j}\left<\left\{\bm{\psi}_{i}\right\},\bm{H}^{\mathrm{inc}}\right>_{\mathrm{P}}\\ \end{bmatrix}\\ &\mathbf{V}^{\mathrm{M}}=\begin{bmatrix}\mathbf{O}\\ \mathrm{j}\mathbf{V}_{\mathrm{P}}^{m}\\ \end{bmatrix},\left[\mathbf{V}_{\mathrm{P}}^{m}\right]_{i}=-2\left<\bm{\psi}_{i},\sum_{\alpha}{v_{\alpha}\frac{\hat{n}\times\mathbf{e}_{\alpha}}{\sqrt{\eta_{\alpha}}}}\right>_{\mathrm{P}}\end{split}$ (16)
@@ V-A Validation example
Figure 5: Gain pattern at 3.5 GHz. (a) xoz-plane, (b) yoz-plane.